### v1

# A Unified Spectral Framework for Aging, Heterogeneous, and Distributed Order Systems via Weighted Weyl-Sonine Operators

Abstract
Standard fractional calculus has successfully modeled systems with power-law memory. However, complex phenomena in heterogeneous media often exhibit multi-scale memory effects and aging properties that classical operators cannot capture. In this work, we construct a unified framework by defining the Weighted Weyl-Sonine Operators. This formalism offers a fundamental generalization of fractional calculus, freeing the theory from the constraints of power-law memory (via Sonine kernels), time-translation invariance (via scale and weight functions), and artificial history truncation (via Weyl integration).
The main result is a Generalized Spectral Mapping Theorem, proving that the Weighted Fourier Transform acts as a universal diagonalization map for these operators. We rigorously characterize the admissible memory kernels through the class of Complete Bernstein Functions, ensuring that the resulting operators preserve the fundamental properties of positivity and monotonicity. Furthermore, we establish a theoretical bridge between the algebraic Sonine definition and the analytical Marchaud representation involving Lévy measures.
Finally, we apply this theory to solve generalized relaxation equations and Weighted Distributed Order evolution problems, demonstrating that phenomena of ultra-slow diffusion and retarded aging can be treated explicitly within this unified spectral framework.
Keywords: Weighted fractional calculus; Sonine kernels; Spectral analysis; Bernstein functions; Distributed order; Aging systems.
MSC 2020: 26A33, 42A38, 47G10.
1 Introduction
The framework of Fractional Calculus has firmly established itself as a fundamental tool for modeling anomalous dynamics in complex systems. Traditionally, operators based on the Riemann-Liouville or Caputo definitions have been the gold standard for describing processes with power-law memory. These operators have successfully captured phenomena in linear viscoelasticity [24, 25], anomalous transport and diffusion [26], and dielectric relaxation processes [28, 27]. The ubiquity of the power-law kernel $k(t)\propto t^{-\alpha}$ arises from its scale-invariance, which naturally models the self-similar fractal structures inherent to these media [29, 30].
However, experimental evidence in highly heterogeneous materials—such as biological tissues, amorphous polymers, and porous media—suggests that memory effects are often far more intricate. In these systems, relaxation rates may evolve over time due to structural changes (“aging”), or the memory function may exhibit transitions between different decay regimes. For instance, in aging polymers or hardening concrete, the system’s response depends explicitly on its age, violating the time-translation invariance inherent to classical convolution operators [17, 18]. Furthermore, complex porous media often exhibit ultra-slow diffusion best described by distributed orders [19], while inertial viscoelasticity can lead to oscillating relaxation dynamics that monotonic power-laws fail to capture [21].
To address these limitations, the theory of General Fractional Calculus (GFC) has emerged as a rigorous generalization, replacing the standard power-law with arbitrary kernels $k(t)$ satisfying the Sonine condition [2]. Yet, a fundamental question remains regarding the domain of these operators. As eloquently argued by Stinga [4], true “memory effects”—likened to an elephant never forgetting its complete history—are best captured by operators of the Weyl-Marchaud type defined on the entire real line $\mathbb{R}$. Restricting analysis to $t>0$ artificially truncates the system’s history, whereas a complete thermodynamic description requires integrating over $(-\infty,t)$.
Recently, significant strides have been made in understanding the structure of these generalized operators. A major breakthrough was achieved by Al-Refai and Fernandez [10], who provided a unified algebraic framework for Sonine operators via the method of conjugations. Their work elegantly demonstrates that many weighted operators can be constructed by conjugating standard derivatives with invertible linear operators. However, their approach implies an algebraic treatment typically using the Laplace transform on semi-axes. Consequently, the spectral analysis of “full history” operators on the unbounded domain $\mathbb{R}$—particularly those combining general memory kernels with time-dependent weights—remains an open problem. While abstract approaches to evolution equations have recently appeared (e.g., Lizama and Murillo-Arcila, 2025 [23]), an explicit harmonic analysis framework allowing for concrete calculations in these heterogeneous settings is still lacking.
This gap has led to a fragmentation of the field into a “zoo” of disconnected operators. Researchers must often choose between the power-law memory of Riemann-Liouville, the logarithmic scaling of Hadamard, the non-singular kernels of Caputo-Fabrizio, or tempered behaviors. While the Generalized Fractional Calculus of Luchko provides the algebraic structure for arbitrary kernels, the physical modeling of aging viscoelasticity, inertial oscillation, and ultra-slow diffusion demands a framework that simultaneously handles general memory geometry and time-domain heterogeneity.
In a recent study [7], we laid the foundations for a unified approach by constructing weighted Weyl derivatives for the specific case of power-law kernels. In this work, we significantly extend those results by constructing the Weighted Weyl-Sonine Operators. This class acts as a “master framework” that generalizes operators of differentiation and integration by incorporating three independent degrees of freedom:
  • The Sonine kernel $k(z)$, which determines the relaxation spectrum (e.g., monotonic vs. oscillatory).
  • The scale function $\psi(t)$, which determines the domain geometry (e.g., invariance of scale vs. translation).
  • The weight function $\omega(t)$, which controls aging or tempering effects.
Our main contribution is a Generalized Spectral Mapping Theorem, which proves that the Weighted Fourier Transform acts as a universal diagonalization map for these complex operators. This result is non-trivial, as it requires handling the interplay between the Sonine convolution condition and non-translation-invariant weight functions.
The paper is organized as follows: Section 2 reviews the preliminaries. Section 3 constructs the operators and proves the Spectral Mapping Theorem. Finally, Section 4 applies this theory to two distinct cases: a generalized relaxation equation with oscillating memory (Bessel kernel) and the problem of distributed-order evolution in aging media, deriving explicit solutions via the spectral method.
2 Preliminaries and Functional Framework
In this section, we summarize the necessary functional spaces and integral transform tools developed in [8], and we recall the fundamental properties of Sonine kernels [2, 9].
2.1 General Fractional Calculus with Sonine Kernels
Following the framework established by Luchko [2, 9], we first review the construction of general fractional operators on the semi-axis $\mathbb{R}_{+}$. This serves as the local foundation for our non-local Weyl operators.
Definition 2.1 (Space $C_{-1}$).
Let $C_{-1}(0,\infty)$ be the space of functions $f$ defined on $(0,\infty)$ such that $f(t)=t^{p}g(t)$ for some real number p>1p>-1 and $g\in C[0,\infty)$. This space contains both functions continuous at the origin (if $p\geq 0$) and functions with integrable singularities (if 1<p<0-1<p<0).
The core of this theory relies on the so-called Sonine condition, which generalizes the relation between the power-law kernels of the Riemann-Liouville calculus.
Definition 2.2 (The Sonine Condition).
Let $k,\kappa\in C_{-1}(0,\infty)$ be two real-valued kernels. The pair $(k,\kappa)$ is said to satisfy the Sonine condition if their Dirichlet convolution is identically unity:
$(k*\kappa)(t)=\int_{0}^{t}k(t-\tau)\kappa(\tau)\,d\tau=1,\quad\forall t>0.$ (1)
The set of all such pairs $\mathcal{K}$ is called the set of Sonine kernels.
Based on these kernels, Luchko defines the General Fractional Integral (GFI) and Derivative (GFD) which unify previous fractional operators.
Definition 2.3 (General Fractional Operators).
For a kernel $k$ satisfying the Sonine condition, the General Fractional Integral denoted by $\mathcal{I}_{(k)}$ is defined as the convolution:
$\mathcal{I}_{(k)}f(t):=(k*f)(t)=\int_{0}^{t}k(t-\tau)f(\tau)\,d\tau,\quad t>0.$ (2)
The associated General Fractional Derivative (GFD) of regularized type, denoted by $\mathbb{D}_{(\kappa)}$, is defined using the associated kernel $\kappa$:
$\mathbb{D}_{(\kappa)}f(t):=\frac{d}{dt}\int_{0}^{t}\kappa(t-\tau)f(\tau)\,d\tau-\kappa(t)f(0).$ (3)
Definition 2.4 (Generalized Sonine Pairs of Arbitrary Order).
Following the framework introduced by Luchko [9], let $n\in\mathbb{N}$ be a natural number. We say that a kernel $\kappa$ belongs to the set of generalized Sonine kernels $\mathcal{L}_{n}$ if there exists an associated kernel $k$ such that the pair $(\kappa,k)$ satisfies the generalized convolution condition:
$(\kappa*k)(t)=\{1\}^{*n}(t)=\frac{t^{n-1}}{(n-1)!},\quad\forall t>0,$ (4)
where $\{1\}^{*n}(t)$ denotes the $n$-fold convolution of the unit function.
Remark 2.5.
For $n=1$, this recovers the classical Sonine condition. For $n>1$, this definition allows for the construction of General Fractional Derivatives of arbitrary order, defined by Luchko as:
$\mathbb{D}_{(k)}f(t)=\frac{d^{n}}{dt^{n}}(k*f)(t)-\sum_{j=0}^{n-1}f^{(j)}(0)\frac{d^{n}}{dt^{n}}(k*\{1\}^{*(j+1)})(t).$ (5)
This formulation is essential for modeling high-order distributed systems where the memory kernel does not necessarily behave as a singular power law near the origin.
Remark 2.6 (Transition to Weighted Weyl Operators).
In our work, we extend these concepts to the entire real line (Weyl type) and introduce time-dependent weights. While Luchko’s operators $\mathbb{D}_{(\kappa)}$ are defined for $t>0$ and involve initial conditions ($f(0)$), our Weighted Weyl operators act on the complete history $t\in(-\infty,\infty)$ and naturally vanish at -$\infty$ for functions in the Schwartz space, removing the need for initial value terms while preserving the algebraic structure of the Sonine pairs.
2.2 Weighted Spaces and Conjugation
Let $\omega\in C^{\infty}(\mathbb{R})$ be a strictly positive weight function and let $\psi\in C^{\infty}(\mathbb{R})$ be a time-scale function satisfying the following conditions:
  1. 1.
    Surjectivity: $\psi(\mathbb{R})=\mathbb{R}$ (covering the full timeline).
  2. 2.
    Non-degeneracy: $\psi^{\prime}(t)>0$ for all $t\in\mathbb{R}$.
The condition $\psi(\mathbb{R})=\mathbb{R}$ ensures that the induced spectral domain covers the entire frequency line $\xi\in\mathbb{R}$, while $\psi^{\prime}(t)>0$ guarantees that the change of variables is a global diffeomorphism, preventing singularities in the Jacobian.
We consider the weighted $L^{2}$-space denoted by $L^{2}_{\psi,\omega}(\mathbb{R})$, equipped with the inner product:
$\langle u,v\rangle_{\psi,\omega}=\int_{-\infty}^{\infty}u(t)\overline{v(t)}\omega(t)\psi^{\prime}(t)\,dt.$ (6)
To handle the heterogeneity of the medium, we utilize the conjugation map approach proposed by Fernandez and Fahad [3]. We define the isometric isomorphism $\mathcal{T}_{\psi,\omega}:L^{2}_{\psi,\omega}(\mathbb{R})\to L^{2}(\mathbb{R})$ as:
$(\mathcal{T}_{\psi,\omega}u)(\xi)=\left(u\circ\psi^{-1}\right)(\xi)\cdot\left(\omega\circ\psi^{-1}\right)(\xi).$ (7)
This operator acts as a "rectifier," mapping functions from the deformed, heterogeneous time-space to the homogeneous standard space.
2.3 The Weighted Fourier Transform
To diagonalize operators with time-dependent coefficients, we employ the Weighted Fourier Transform constructed in our previous work [6]. While the general theory was developed for $\mathbb{R}^{n}$, here we present the one-dimensional formulation adapted to the global time domain $\mathbb{R}$.
We introduce the weighted Lebesgue spaces, denoted by $L^{p}_{\psi,\omega}(\mathbb{R})$, defined as the completion of $C_{c}^{\infty}(\mathbb{R})$ with respect to the norm associated with the weighted measure $d\mu_{\psi,\omega}(t)=\omega(t)\psi^{\prime}(t)dt$:
$\|u\|_{L^{p}_{\psi,\omega}}:=\left(\int_{\mathbb{R}}|u(t)|^{p}\omega(t)\psi^{\prime}(t)\,dt\right)^{1/p}<\infty.$ (8)
Definition 2.7.
For a function $f\in L_{\psi,\omega}^{1}(\mathbb{R})$, the Weighted Fourier Transform with respect to the scale $\psi$ and weight $\omega$, denoted by $\mathcal{F}_{\psi,\omega}$, is defined as:
$\mathcal{F}_{\psi,\omega}f](\xi)=\frac{1}{\sqrt{2\pi}}\int_{-\infty}^{\infty}e^{-i\xi\psi(t)}\omega(t)f(t)\psi^{\prime}(t)\,dt.$ (9)
Analytically, this operator acts as a conjugation of the classical Fourier transform $\mathcal{F}$ by structural isomorphisms. Specifically, if we define the composition operator $Q_{\psi}g(t)=g(\psi(t))$ and the multiplication operator $M_{\omega}f(t)=\omega(t)f(t)$, the definition (9) admits the factorization:
$\mathcal{F}_{\psi,\omega}=\mathcal{F}\circ Q_{\psi}^{-1}\circ M_{\omega}.$ (10)
This factorization is crucial, as it transfers the unitarity and inversion properties of the classical transform directly to the weighted setting via the induced isometry. The inverse transform allows for the recovery of the time-domain solution.
Theorem 2.8 (Inversion Formula).
Let $f$ be a function such that both $f$ and $\mathcal{F}_{\psi,\omega}f$ are integrable in the appropriate weighted sense. Then, the inverse transform is given pointwise by:
$f(t)=\frac{1}{\omega(t)\sqrt{2\pi}}\int_{-\infty}^{\infty}e^{i\xi\psi(t)}[\mathcal{F}_{\psi,\omega}f](\xi)\,d\xi.$ (11)
Proof.
The proof follows directly from the $n$-dimensional case presented in [6] by setting $n=1$, $\phi(x)=\psi(t)$, and noting that the Jacobian determinant $|J_{\phi}(x)|$ reduces to $\psi^{\prime}(t)$ since $\psi$ is strictly increasing. By the change of variables $u=\psi(t)$, the integral reduces to the classical Fourier inversion theorem. ∎
Remark 2.9.
A fundamental property of $\mathcal{F}_{\psi,\omega}$ is that it extends to a unitary operator (isometry) on $L^{2}_{\psi,\omega}(\mathbb{R})$ under suitable conditions on $\omega$. This generalized Plancherel identity (see [6]) is essential for the energy estimates derived in Section 4.
Definition 2.10 (Weighted Schwartz Space as a Fréchet Space).
Let $\psi:\mathbb{R}\to\mathbb{R}$ be a smooth diffeomorphism and let $\omega\in C^{\infty}(\mathbb{R})$ be a non-vanishing weight. The Weighted Schwartz Space, denoted by $\mathcal{S}_{\psi,\omega}(\mathbb{R})$, is defined as the isomorphic image of the standard Schwartz space $\mathcal{S}(\mathbb{R})$ under the mapping $\mathcal{T}_{\psi,\omega}=M_{\omega}^{-1}\circ Q_{\psi}$:
$\mathcal{S}_{\psi,\omega}(\mathbb{R}):=\left\{u:\mathbb{R}\to\mathbb{C}\mid\exists f\in\mathcal{S}(\mathbb{R})\text{ such that }u(t)=\frac{f(\psi(t))}{\omega(t)}\right\}.$ (12)
We equip $\mathcal{S}_{\psi,\omega}(\mathbb{R})$ with the locally convex topology induced by this bijection. Specifically, convergence $u_{n}\to u$ in $\mathcal{S}_{\psi,\omega}$ corresponds to the convergence of the pull-backs $f_{n}=\omega\cdot(u_{n}\circ\psi^{-1})$ in the standard Fréchet topology of $\mathcal{S}(\mathbb{R})$.
Proposition 2.11 (Density and Boundary Decay).
The space $\mathcal{S}_{\psi,\omega}(\mathbb{R})$ is dense in the weighted Hilbert space $L^{2}_{\psi,\omega}(\mathbb{R})$. Furthermore, the rapid decay inherited from $\mathcal{S}(\mathbb{R})$ ensures that for any $u\in\mathcal{S}_{\psi,\omega}(\mathbb{R})$, both the function and its weighted derivatives vanish at infinity:
$\lim_{t\to\pm\infty}u(t)=0\quad\text{and}\quad\lim_{t\to\pm\infty}\mathfrak{D}_{\psi,\omega}^{n}u(t)=0\quad\forall n\in\mathbb{N},$ (13)
where $\mathfrak{D}_{\psi,\omega}^{n}$ denotes the iterated weighted differential operator defined by:
$\mathfrak{D}_{\psi,\omega}^{n}u(t):=\frac{1}{\omega(t)}\left(\frac{1}{\psi^{\prime}(t)}\frac{d}{dt}\right)^{n}\big(\omega(t)u(t)\big).$ (14)
Proof.
The proof relies on the unitary equivalence established by the mapping $\mathcal{T}_{\psi,\omega}$.
1. Density: Recall that the standard Schwartz space $\mathcal{S}(\mathbb{R})$ is dense in $L^{2}(\mathbb{R})$. Since the operator $\mathcal{T}_{\psi,\omega}=M_{\omega}^{-1}Q_{\psi}$ is a unitary isomorphism from $L^{2}(\mathbb{R})$ to $L^{2}_{\psi,\omega}(\mathbb{R})$ (preserving the inner product structure), it maps dense sets to dense sets. Therefore, the image $\mathcal{S}_{\psi,\omega}(\mathbb{R})=\mathcal{T}_{\psi,\omega}(\mathcal{S}(\mathbb{R}))$ is necessarily dense in $L^{2}_{\psi,\omega}(\mathbb{R})$.
2. Boundary Decay: Let $u\in\mathcal{S}_{\psi,\omega}(\mathbb{R})$. By definition, there exists a unique $f\in\mathcal{S}(\mathbb{R})$ such that $u(t)=f(\psi(t))/\omega(t)$, or equivalently $\omega(t)u(t)=f(\psi(t))$. Substituting this into the definition of the weighted derivative (14), we apply the chain rule:
$\mathfrak{D}_{\psi,\omega}^{n}u(t)=\frac{1}{\omega(t)}\left(\frac{1}{\psi^{\prime}(t)}\frac{d}{dt}\right)^{n}f(\psi(t)).$
Let $y=\psi(t)$. Since $\frac{1}{\psi^{\prime}(t)}\frac{d}{dt}=\frac{d}{dy}$, the operator simplifies strictly to the classical derivative in the transformed domain:
$\left(\frac{1}{\psi^{\prime}(t)}\frac{d}{dt}\right)^{n}f(\psi(t))=f^{(n)}(y)\Big|_{y=\psi(t)}.$
Thus, we have the relation:
$\omega(t)\cdot\mathfrak{D}_{\psi,\omega}^{n}u(t)=f^{(n)}(\psi(t)).$
Since $f\in\mathcal{S}(\mathbb{R})$, its derivatives $f^{(n)}$ decay rapidly at infinity. Assuming the weight $\omega(t)$ does not grow superexponentially relative to the decay of $f$ (a condition satisfied by admissible weights), the term on the RHS vanishes as $\psi(t)\to\pm\infty$. Consequently, $\mathfrak{D}_{\psi,\omega}^{n}u(t)\to 0$ as $t\to\pm\infty$. ∎
3 The Weighted Weyl-Sonine Framework and Spectral Analysis
In this section, we construct the unified operator. To ensure the convergence of integrals over the history $(-\infty,t]$, we must adapt the classical Sonine conditions (typically defined on bounded intervals or $\mathbb{R}_{+}$) to the context of Weyl operators on the entire real line.
3.1 Admissible Kernels and Operators
Unlike the Riemann-Liouville case, where integration starts at a finite point $a$, the Weyl operator integrates the entire history. This requires the memory kernel to decay sufficiently fast as time delay increases.
Definition 3.1 (Generalized Tempered Sonine-Weyl Pair of Order $n$).
Let $n\in\mathbb{N}$. Let $k,\kappa\in\mathcal{S}^{\prime}(\mathbb{R})$ be two tempered distributions. The pair $(k,\kappa)$ is called an admissible tempered pair of order $n$ if the following conditions are satisfied:
  1. 1.
    Support Condition: Both distributions are supported on the closed semi-axis $0,\infty)$. That is, $k,\kappa\in\mathcal{S}^{\prime}_{+}(\mathbb{R})$.
  2. 2.
    Distributional Generalized Condition: The convolution of the kernels yields the $n$-th primitive of the Dirac delta, which corresponds to the truncated power function distribution:
    $k*\kappa=\frac{t_{+}^{n-1}}{(n-1)!}\quad\text{in }\mathcal{S}^{\prime}(\mathbb{R}),$ (15)
    where $t_{+}^{\lambda}$ is the standard truncated power distribution defined by $\langle t_{+}^{\lambda},\phi\rangle=\int_{0}^{\infty}t^{\lambda}\phi(t)dt$.
    For the classical case $n=1$, this recovers the Heaviside step function: $k*\kappa=\theta(t)$.
  3. 3.
    Spectral Condition: To ensure the operator corresponds to a fractional process, the Laplace symbol $\hat{k}(s)$ (defined for $\text{Re}(s)>0$) satisfies algebraic growth/decay conditions at infinity consistent with a fractional order strictly between n1n-1 and $n$.
Remark 3.2 (From Local Functions to Global Distributions).
Note that any classical Sonine kernel $k_{loc}(t)\in C_{-1}(0,\infty)$ identified in Section 2.1 generates a regular tempered distribution $k\in\mathcal{S}^{\prime}(\mathbb{R})$ by extending it as zero for $t<0$. Condition (15) generalizes the classical relation $\int_{0}^{t}k(t-\tau)\kappa(\tau)d\tau=1$ (case $n=1$), allowing us to include kernels that involve singular distributions (like $\delta(t)$) which are essential for modeling elastic solids or purely viscous fluids, but are excluded in the classical function-space theory.
Remark 3.3 (Consistency with Classical and Generalized Kernels).
This definition unifies the classical and generalized frameworks within distribution theory:
  • Case $n=1$ (Classical): We recover the standard Sonine condition $k*\kappa=1$ (for $t>0$). This covers standard models like subdiffusion, where kernels are singular at the origin.
  • Case $n>1$ (High-Order): This allows for kernels that are continuous at the origin (regular kernels), as discussed in Remark LABEL:rem:singularity_discussion. In this setting, the operator $\mathfrak{D}_{\psi,\omega}^{(\kappa)}$ acts as a weighted derivative of order $n$, involving the $n$-th derivative of the convolution.
We now introduce the main operator of this work, constructed via the conjugation of the classical Sonine integral with the weight and scale operators defined in [3].
Definition 3.4 (Auxiliary Structure Operators).
Let $(\psi,\omega)$ be an admissible structure pair. We define the following operators acting on measurable functions $f:\mathbb{R}\to\mathbb{C}$:
  1. 1.
    The Multiplication Operator $M_{\omega}$, which weights the function:
    $(M_{\omega}f)(t):=\omega(t)f(t).$ (16)
  2. 2.
    The Substitution (Time-Warping) Operator $Q_{\psi}$, which deforms the domain:
    $(Q_{\psi}f)(t):=f(\psi(t)).$ (17)
Since $\psi$ is a diffeomorphism and $\omega$ is non-vanishing, both operators are invertible on suitable spaces. Specifically, the inverse substitution is given by $(Q_{\psi}^{-1}g)(t)=g(\psi^{-1}(t))$.
Remark 3.5.
These operators allow us to factorize the Weighted Fourier Transform as a composition: $\mathcal{F}_{\psi,\omega}=\mathcal{F}\circ Q_{\psi}^{-1}\circ M_{\omega}$. This factorization is the key to the spectral analysis.
Definition 3.6 (Left-Sided Weighted Weyl-Sonine Integral).
Let $(k,\kappa)$ be an admissible Sonine pair and let $u$ be a function in the weighted space $\mathcal{S}_{\psi,\omega}$. We define the Left-Sided Weighted Weyl-Sonine Integral, denoted by $\mathfrak{I}_{\psi,\omega}^{(k)}$, as follows:
1. Explicit Form: The operator is defined pointwise by the weighted integral:
$\mathfrak{I}_{\psi,\omega}^{(k)}u(t):=\frac{1}{\omega(t)}\int_{-\infty}^{t}k\Big(\psi(t)-\psi(\tau)\Big)\omega(\tau)u(\tau)\psi^{\prime}(\tau)\,d\tau.$ (18)
2. Structural Form: In terms of the auxiliary operators from Definition 3.4, the integral acts as a conjugated convolution:
$\mathfrak{I}_{\psi,\omega}^{(k)}=M_{\omega}^{-1}Q_{\psi}(k*)Q_{\psi}^{-1}M_{\omega},$ (19)
where (k)(k*) denotes the standard convolution operator on $\mathbb{R}$.
Physical Interpretation: This operator models a process with memory in a heterogeneous medium:
  • The term $\Delta_{\psi}(t,\tau)=\psi(t)-\psi(\tau)$ represents the subjective time elapsed between the past event $\tau$ and the present $t$.
  • The kernel $k(\Delta_{\psi})$ assigns a weight to this memory (typically singular at the origin).
  • The ratio $\omega(\tau)/\omega(t)$ accounts for the aging or changing density of the medium during the process.
Definition 3.7 (Weighted Weyl-Sonine Derivative of Order $n$).
Let $n\in\mathbb{N}$ and let $(k,\kappa)$ be an admissible tempered Sonine-Weyl pair of order $n$. We define the Weighted Weyl-Sonine derivative, denoted by $\mathfrak{D}_{\psi,\omega}^{(\kappa)}$, as the regularized inverse of the integral operator.
1. Explicit Form: For a suitable function $u$, the operator is defined pointwise by applying the weighted differential operator $n$ times:
$\mathfrak{D}_{\psi,\omega}^{(\kappa)}u(t)=\frac{1}{\omega(t)}\left(\frac{1}{\psi^{\prime}(t)}\frac{d}{dt}\right)^{n}\int_{-\infty}^{t}\kappa(\psi(t)-\psi(\tau))\omega(\tau)u(\tau)\psi^{\prime}(\tau)d\tau.$ (20)
2. Structural Form: By identifying the differential operator $\frac{1}{\psi^{\prime}(t)}\frac{d}{dt}$ with the derivative $\frac{d}{dy}$ in the warped time $y=\psi(t)$, the operator admits the following conjugation structure:
$\mathfrak{D}_{\psi,\omega}^{(\kappa)}=M_{\omega}^{-1}Q_{\psi}\left(\frac{d^{n}}{dy^{n}}\circ(\kappa*)\right)Q_{\psi}^{-1}M_{\omega}.$ (21)
This abstract form emphasizes that our operator is structurally equivalent to a classical convolution derivative of order $n$, conjugated by the geometry $\psi$ and weight $\omega$.
$\mathfrak{D}_{\psi,\omega}^{(\kappa)}=M_{\omega}^{-1}Q_{\psi}\left(\frac{d^{n}}{dy^{n}}\circ(\kappa*)\right)Q_{\psi}^{-1}M_{\omega}.$ (22)
This abstract form emphasizes that our operator is structurally equivalent to a classical convolution derivative of order $n$, conjugated by the geometry $\psi$ and weight $\omega$.
Note that for $n=1$, the operator simplifies to the weighted first derivative $\frac{1}{\psi^{\prime}(t)}\frac{d}{dt}$, recovering the classical case. For $n\geq 2$, this formulation captures the higher-order dynamics required by the generalized Sonine pairs.
Remark 3.8 (Algebraic Structure and the Global Equivalence of Forms).
The construction in Definitions 3.6-3.7 hereditarily preserves the convolution algebra of the standard Sonine calculus, albeit non-trivially deformed by the diffeomorphism $\psi$ and the weight $\omega$.
This global structure also resolves the distinction between Riemann-Liouville and Caputo definitions commonly found in General Fractional Derivatives on finite intervals $[0,T]$ (e.g., Luchko [9]). In those local settings, the non-commutativity of differentiation and integration necessitates explicit regularization terms (initial values). In contrast, the rapid decay of test functions in our weighted Schwartz space $\mathcal{S}_{\psi,\omega}(\mathbb{R})$ ensures that all boundary terms vanish naturally:
$\lim_{t\to-\infty}\left(\frac{1}{\psi^{\prime}(t)}\frac{d}{dt}\right)^{k}u(t)=0,\quad\forall k<n.$
Consequently, the differential and integral operators commute strictly over the domain, making the Riemann-Liouville and Caputo forms mathematically equivalent. While physically interpreted as a "hypothesis of quiescent past," analytically this is a direct consequence of the topological structure of the functional space.
Remark 3.9 (Unification of Fractional Calculi).
A major advantage of the proposed framework is its ability to encompass a vast array of existing fractional operators on the real line as particular cases. By carefully selecting the memory kernel $k(z)$, the scale function $\psi(t)$, and the weight $\omega(t)$, the operator $\mathfrak{D}_{\psi,\omega}^{(k)}$ reduces to the operators listed in Table 1.
Table 1: Special cases of the Weighted Weyl-Sonine Operator $\mathfrak{D}_{\psi,\omega}^{(k)}$ acting on $\mathbb{R}$ (or valid sub-domains).
Operator Name Kernel $k(z)$ Scale $\psi(t)$ Weight $\omega(t)$
Standard Weyl $\frac{z^{-\alpha}}{\Gamma(1-\alpha)}$ $t$ 11
Tempered Weyl $\frac{z^{-\alpha}}{\Gamma(1-\alpha)}$ $t$ $e^{\lambda t}$
Hadamard-type $\frac{z^{-\alpha}}{\Gamma(1-\alpha)}$ $\ln t$ 11
Katugampola-type $\frac{z^{-\alpha}}{\Gamma(1-\alpha)}$ $t^{\rho}$ 11
Caputo-Fabrizio $\exp\left(-\frac{\alpha}{1-\alpha}z\right)$ $t$ 11
Atangana-Baleanu $E_{\alpha}\left(-\frac{\alpha}{1-\alpha}z^{\alpha}\right)$ $t$ 11
Prabhakar (Generalized) $z^{\beta-1}E_{\rho,\beta}^{\gamma}(-\lambda z^{\rho})$ $t$ 11
Distributed Order $\int_{0}^{1}b(\alpha)\frac{z^{-\alpha}}{\Gamma(1-\alpha)}d\alpha$ $t$ 11
Aging Viscoelasticity $\frac{z^{-\alpha}}{\Gamma(1-\alpha)}$ $\psi(t)$ $\omega(t)$
Table 1 illustrates the unifying capability of the proposed framework. While the first columns encompass geometric variations (time-scales and kernel shapes), the inclusion of the weight $\omega(t)$ reveals a deeper algebraic connection between weighted spaces and tempered operators. In particular, the case designated as ‘Tempered Weyl’ ($\omega(t)=e^{\lambda t}$) is often conceptually distinguished from ‘Tempered Memory’ models (where the kernel itself decays as $e^{-\lambda t}k(t)$). However, our formalism reveals that these two approaches are algebraically equivalent.
Remark 3.10 (Equivalence of Space Tempering and Memory Tempering).
It is a non-trivial fact that weighting the function space is isomorphic to tempering the memory kernel. Consider the weighted integral $\mathfrak{I}_{\psi,\omega}^{(k)}$ with standard scale $\psi(t)=t$ and exponential weight $\omega(t)=e^{\lambda t}$. A direct computation shows:
$\mathfrak{I}_{t,e^{\lambda t}}^{(k)}u(t)=e^{-\lambda t}\int_{-\infty}^{t}k(t-\tau)e^{\lambda\tau}u(\tau)\,d\tau=\int_{-\infty}^{t}\left[e^{-\lambda(t-\tau)}k(t-\tau)\right]u(\tau)\,d\tau.$ (23)
The term in brackets identifies a new effective kernel $k_{\lambda}(z)=e^{-\lambda z}k(z)$. Thus, the weighted operator on standard $L^{2}$ functions is identical to the unweighted convolution operator with a tempered kernel. This confirms that our framework naturally generates the entire class of tempered fractional operators (such as the Tempered RL or Caputo derivatives) simply by selecting the appropriate weight $\omega$, without needing to redefine the kernel class ad-hoc.
3.2 Spectral Analysis
To establish a direct connection with the classical theory of pseudo-differential operators, we characterize the Generalized Sonine pair through its spectral signature.
Remark 3.11 (The Spectral Advantage).
Luchko notes in [2] that an analytical description of the entire set of Sonine kernels remains an open problem. In the time domain, explicit closed-form formulas for the pairs $(k,\kappa)$ are often intractable or involve special functions that are computationally expensive to evaluate (e.g., H-functions).
Our spectral framework bypasses this hurdle. By characterizing the operator via its symbol $\Phi(s)$ rather than its temporal kernel, we can solve fractional differential equations analytically even when the explicit form of the memory kernel $\kappa(t)$ is unknown, provided its Laplace symbol is well-defined. This is particularly useful in data-driven modeling, where $\Phi(s)$ is often fitted directly from rheological spectra.
Definition 3.12.
[The Spectral Symbol of Order $n$] Let $n\in\mathbb{N}$ and let $(k,\kappa)$ be an admissible Sonine-Weyl pair of order $n$. We define the Spectral Symbol $\Phi(s)$ (for $\text{Re}(s)>0$) as the Laplace transform of the associated derivative kernel $\kappa$:
$\Phi(s):=\mathcal{L}\{\kappa\}(s).$ (24)
Recalling from Definition 3.1 that $(k*\kappa)(t)=t_{+}^{n-1}/(n-1)!$, the Laplace transform of the convolution implies the fundamental algebraic constraint:
$\mathcal{L}\{k\}(s)\cdot\Phi(s)=\frac{1}{s^{n}}.$ (25)
Consequently, the integral kernel $k$ is fully determined in the spectral domain by the inverse symbol:
$\mathcal{L}\{k\}(s)=\frac{1}{s^{n}\Phi(s)}.$ (26)
Proposition 3.13 (Existence and Analytic Structure of Admissible Pairs).
Let $\Phi(s)$ be a function analytic in the right half-plane. A unique pair of kernels $(k,\kappa)$ satisfying the Sonine relation exists if and only if they are determined by the inverse Laplace transforms of (27):
$\mathcal{L}\{k\}(s)=\frac{1}{s^{n}\Phi(s)},\quad\mathcal{L}\{\kappa\}(s)=\Phi(s).$ (27)
Moreover, to ensure physical realizability (non-negativity and passivity), if $\Phi(s)$ belongs to the class of Complete Bernstein Functions (CBF), then:
  1. 1.
    The derivative kernel $\kappa$ is a tempered distribution representing a Lévy measure (possibly with a drift), ensuring the operator models a sub-diffusive or relaxational process.
  2. 2.
    The integral kernel $k(t)$ is a locally integrable, completely monotone function on $(0,\infty)$, i.e.,
    $(-1)^{j}k^{(j)}(t)\geq 0$ for all $j\in\mathbb{N}_{0}$.
Proof.
Algebraic Existence. The convolution condition $(k*\kappa)(t)=t_{+}^{n-1}/(n-1)!$ is strictly equivalent in the Laplace domain to the algebraic equation $\hat{k}(s)\hat{\kappa}(s)=s^{-n}$. Identifying $\hat{\kappa}(s)=\Phi(s)$ immediately yields the unique solution for $\hat{k}(s)$ given in (26).
Analytic Properties. Assume $\Phi\in\mathcal{CBF}$. By the characterization of Complete Bernstein Functions (see Schilling et al. [20], Thm 6.2), the reciprocal $1/\Phi(s)$ is a Stieltjes function. Stieltjes functions $f(s)$ possess the integral representation:
$f(s)=\frac{a}{s}+b+\int_{0}^{\infty}\frac{1}{s+r}\,d\sigma(r),\quad a,b\geq 0,\,\sigma\geq 0.$
It is a classical result that for any Stieltjes function $f(s)$, the product $s^{-n}f(s)$ (for $n\geq 1$) remains the Laplace transform of a completely monotone function. Thus, the kernel $k(t)=\mathcal{L}^{-1}\{s^{-n}/\Phi(s)\}$ satisfies $(-1)^{j}k^{(j)}(t)\geq 0$.
This monotonicity ensures that the integral operator $\mathfrak{I}_{\psi,\omega}^{(k)}$ is positivity-preserving, a necessary condition for the validity of maximum principles and the well-posedness of the associated Cauchy problem. ∎
We now proceed to the main spectral mapping theorem. It is worth noting that the classical fractional calculus algebra—characterized by the power law $s^{-\alpha}$—is recovered here as a particular instance, now generalized by the spectral symbol $\Phi(s)$.
Theorem 3.14 (Generalized Spectral Mapping Theorem).
Let $n\in\mathbb{N}$ and let $u\in\mathcal{S}_{\psi,\omega}(\mathbb{R})$. Let $(k,\kappa)$ be an admissible Sonine pair with associated Laplace symbol $\Phi(s):=\hat{\kappa}(s)$. The Weighted Weyl-Sonine operators satisfy the following spectral relations:
  1. 1.
    For the Integral:
    $\mathcal{F}_{\psi,\omega}\{\mathfrak{I}_{\psi,\omega}^{(k)}u\}(\xi)=\frac{1}{(i\xi)^{n}\Phi(i\xi)}\cdot\mathcal{F}_{\psi,\omega}\{u\}(\xi).$ (28)
  2. 2.
    For the Derivative: Considering the operator $\mathfrak{D}_{\psi,\omega}^{(\kappa)}$ from Definition 3.7, the spectral mapping holds:
    $\mathcal{F}_{\psi,\omega}[\mathfrak{D}_{\psi,\omega}^{(\kappa)}u](\xi)=(i\xi)^{n}\Phi(i\xi)\cdot\mathcal{F}_{\psi,\omega}[u](\xi).$ (29)
Proof.
We rely on the structural factorization of the Weighted Fourier Transform established in Section 3.2: $\mathcal{F}_{\psi,\omega}=\mathcal{F}\circ Q_{\psi}^{-1}\circ M_{\omega}$, which acts on the auxiliary function $v=Q_{\psi}^{-1}(M_{\omega}u)\in\mathcal{S}(\mathbb{R})$.
Case 1 (Integral): Recall from Definition 3.6 that the integral operator admits the conjugation structure:
$\mathfrak{I}_{\psi,\omega}^{(k)}=M_{\omega}^{-1}Q_{\psi}(k*)Q_{\psi}^{-1}M_{\omega}.$
Applying the Weighted Fourier Transform to the operator $\mathfrak{I}_{\psi,\omega}^{(k)}u$ leads to a direct cancellation of the structural maps:
$\mathcal{F}_{\psi,\omega}\left(\mathfrak{I}_{\psi,\omega}^{(k)}u\right)$ $=(\mathcal{F}\circ Q_{\psi}^{-1}\circ M_{\omega})\circ(M_{\omega}^{-1}Q_{\psi}(k*)Q_{\psi}^{-1}M_{\omega}u)$
$=\mathcal{F}\left(k*(Q_{\psi}^{-1}M_{\omega}u)\right)=\mathcal{F}\{k*v\}.$
Since $v\in\mathcal{S}(\mathbb{R})$, we apply the classical Convolution Theorem. Using the Sonine condition $\hat{k}(s)\hat{\kappa}(s)=s^{-n}$, we relate the symbol of $k$ to $\Phi$:
$\mathcal{F}\{k*v\}(\xi)=\hat{k}(i\xi)\cdot\mathcal{F}\{v\}(\xi)=\frac{1}{(i\xi)^{n}\hat{\kappa}(i\xi)}\mathcal{F}\{v\}(\xi).$
Substituting $\Phi(i\xi)=\hat{\kappa}(i\xi)$ and identifying $\mathcal{F}\{v\}=\mathcal{F}_{\psi,\omega}\{u\}$, the first result follows.
Case 2 (Derivative): For the derivative operator of order $n$, we utilize the structural form given in Eq. (22):
$\mathfrak{D}_{\psi,\omega}^{(\kappa)}=M_{\omega}^{-1}Q_{\psi}\left(\frac{d^{n}}{dy^{n}}\circ(\kappa*)\right)Q_{\psi}^{-1}M_{\omega}.$
Applying $\mathcal{F}_{\psi,\omega}$ again cancels the outer geometrical factors, yielding the Fourier transform of a convoluted derivative in the Schwartz space:
$\mathcal{F}_{\psi,\omega}[\mathfrak{D}_{\psi,\omega}^{(\kappa)}u](\xi)=\mathcal{F}\left[\frac{d^{n}}{dy^{n}}(\kappa*v)\right](\xi).$
Using the derivative property of the classical Fourier transform iteratively, $\mathcal{F}\{f^{(n)}\}(\xi)=(i\xi)^{n}\mathcal{F}\{f\}(\xi)$, we obtain:
$\mathcal{F}\left[\frac{d^{n}}{dy^{n}}(\kappa*v)\right](\xi)=(i\xi)^{n}\cdot\mathcal{F}\{\kappa*v\}(\xi)=(i\xi)^{n}\Phi(i\xi)\mathcal{F}\{v\}(\xi).$
Recovering $\mathcal{F}_{\psi,\omega}[u](\xi)$ from $\mathcal{F}\{v\}(\xi)$ completes the proof. ∎
Remark 3.15 (Natural Domains and Future Outlook).
The Generalized Spectral Mapping Theorem implies that the natural energy spaces for these operators are the Weighted Sonine-Sobolev spaces, defined via the spectral magnitude of the symbol $\Phi$:
$\mathcal{H}_{\psi,\omega}^{\Phi}(\mathbb{R}):=\left\{u\in L^{2}_{\psi,\omega}(\mathbb{R}):\int_{\mathbb{R}}\left(1+|\Phi(i\xi)|^{2}\right)|\mathcal{F}_{\psi,\omega}u(\xi)|^{2}\,d\xi<\infty\right\}.$
While the construction of these spaces is naturally suggested by our spectral framework, a rigorous treatment—including density results, embedding theorems, and trace properties involving the weights $(\psi,\omega)$—lies outside the scope of the present work. These analytic aspects constitute a substantial topic that will be addressed in a forthcoming study dedicated to the regularity of solutions.
3.3 The Generalized Marchaud Representation
To fully align our framework with the analytical philosophy of non-local operators (see Stinga [4] and Samko [1]), we establish a connection with hypersingular integral representations. While the Sonine form (convolution) is convenient for algebraic manipulation, the Marchaud form (finite differences) highlights the non-local nature of the operator as an accumulation of ‘jumps’ governed by a Lévy measure.
We show that our Weighted Weyl-Sonine derivative admits such a representation, provided the spectral symbol $\Phi(s)$ corresponds to a Bernstein function associated with differentiation orders strictly between 0 and 1.
Definition 3.16 (Associated Lévy Density).
Let $\Phi(s)$ be the spectral symbol defined in Definition 3.12. Assume $\Phi$ admits the Lévy-Khintchine representation for a drift-free subordinator:
$\Phi(s)=\int_{0}^{\infty}(1-e^{-sz})\,d\mu(z),$ (30)
where $\mu$ is a Lévy measure on $(0,\infty)$ satisfying $\int_{0}^{\infty}\min(1,z)\,d\mu(z)<\infty$. Furthermore, we assume $\mu$ is absolutely continuous with respect to the Lebesgue measure, admitting a Lévy density $g(z)$ such that $d\mu(z)=g(z)dz$.
Theorem 3.17 (Weighted Marchaud Formula).
Under the assumptions of Definition 3.16, the Weighted Weyl-Sonine derivative $\mathfrak{D}_{\psi,\omega}^{(\kappa)}$ admits the following generalized Marchaud representation for any $u\in\mathcal{S}_{\psi,\omega}$:
$\mathfrak{D}_{\psi,\omega}^{(\kappa)}u(t)=\frac{1}{\omega(t)}\int_{-\infty}^{t}\Big((\omega u)(t)-(\omega u)(\tau)\Big)g\Big(\psi(t)-\psi(\tau)\Big)\psi^{\prime}(\tau)\,d\tau.$ (31)
Proof.
We employ the uniqueness of the spectral representation established in Theorem 3.14. Our goal is to show that the right-hand side of (31), under the conjugation map $\mathcal{F}_{\psi,\omega}$, yields the correct spectral symbol $\Phi$.
Let $u\in\mathcal{S}_{\psi,\omega}$ and define the auxiliary function $v=Q_{\psi}^{-1}M_{\omega}u$ on the rectified timeline $y=\psi(t)$. Recall that $(M_{\omega}u)(t)=v(\psi(t))$. Substituting this into the integral term of (31), the difference term becomes:
$(\omega u)(t)-(\omega u)(\tau)=v(\psi(t))-v(\psi(\tau)).$
Let $I(t)$ denote the integral in (31). We analyze the operator on the rectified domain by applying the conjugation steps and evaluating at $t=\psi^{-1}(y)$:
$(Q_{\psi}^{-1}M_{\omega}I)(y)=\int_{-\infty}^{\psi^{-1}(y)}\Big(v(y)-v(\psi(\tau))\Big)g\Big(y-\psi(\tau)\Big)\psi^{\prime}(\tau)\,d\tau.$
We perform the change of variables $z=y-\psi(\tau)$. This implies $dz=-\psi^{\prime}(\tau)d\tau$. The integration limits transform as follows: as $\tau\to$-$\infty$, $z\to\infty$; and as $\tau\to\psi^{-1}(y)$, $z\to 0$. Absorbing the minus sign into the limits, we obtain:
$(Q_{\psi}^{-1}M_{\omega}I)(y)=\int_{0}^{\infty}\Big(v(y)-v(y-z)\Big)g(z)\,dz.$
This represents a convolution of differences. We take the classical Fourier transform $\mathcal{F}$ with respect to $y$. Using the linearity of the integral and Fubini’s theorem (justified by the convergence of the Lévy measure against bounded differences), we interchange the Fourier integral with the $dz$ integral:
$\mathcal{F}\left\{\int_{0}^{\infty}(v(\cdot)-v(\cdot-z))g(z)\,dz\right\}(\xi)$ $=\int_{0}^{\infty}\mathcal{F}\Big\{v(y)-v(y-z)\Big\}(\xi)\,g(z)\,dz$
$=\int_{0}^{\infty}\Big(\hat{v}(\xi)-e^{-i\xi z}\hat{v}(\xi)\Big)g(z)\,dz$
$=\hat{v}(\xi)\left(\int_{0}^{\infty}(1-e^{-i\xi z})g(z)\,dz\right).$
By the Lévy-Khintchine formula (30) (with $s=i\xi$), the integral in parentheses is exactly $\Phi(i\xi)$. Thus, the spectral symbol matches Definition 3.12, completing the proof. ∎
Remark 3.18.
This result unifies the algebraic approach with the analytical perspective. For the classical fractional case where $\Phi(s)=s^{\alpha}$ (with $\alpha\in(0,1)$), the associated Lévy density is $g(z)=\frac{\alpha}{\Gamma(1-\alpha)}z^{-(1+\alpha)}$. Substituting this into (31) recovers the weighted extension of the classical Marchaud fractional derivative.
4 Applications to Anomalous Relaxation and Diffusion
In this final section, we demonstrate the versatility of the Weighted Weyl-Sonine framework by solving two classes of problems that arise in complex media: relaxation equations with oscillating memory and distributed-order evolution equations.
Remark 4.1.
While the theoretical framework developed in Section 3 supports operators of arbitrary order $n$, in this section we focus on the case $n=1$. This covers the vast majority of physical relaxation and diffusion phenomena, where the fundamental dynamics are governed by first-order evolution equations perturbed by memory.
4.1 Spectral Solvability of Generalized Relaxation Models
We consider the fundamental evolution equation governing the relaxation of a field $u\in\mathcal{S}_{\psi,\omega}$ (e.g., stress in a viscoelastic material or density in a porous medium) under the action of the generalized derivative.
Problem Statement: Find a unique solution $u(t)$ to the Cauchy-type problem on the real line:
$\mathfrak{D}_{\psi,\omega}^{(\kappa)}u(t)+\lambda u(t)=f(t),\quad t\in\mathbb{R},$ (32)
where $\lambda>0$ is a relaxation parameter, $f\in L^{2}_{\psi,\omega}(\mathbb{R})$ is a source term, and $\mathfrak{D}_{\psi,\omega}^{(\kappa)}$ is the Weighted Weyl-Sonine derivative (of order $n=1$) associated with the kernel $\kappa$.
Using the Generalized Spectral Mapping Theorem (Theorem 3.14), we can characterize the solvability purely in terms of the spectral symbol $\Phi$.
Theorem 4.2 (Well-posedness and Spectral Resolution).
Let $\Phi(s)=\mathcal{L}\{\kappa\}(s)$ be the spectral symbol of the derivative. Assume the following Ellipticity Condition holds on the imaginary axis:
$\inf_{\xi\in\mathbb{R}}\left|i\xi\Phi(i\xi)+\lambda\right|>0.$ (33)
Then, the operator $(\mathfrak{D}_{\psi,\omega}^{(\kappa)}+\lambda I)$ is a linear isomorphism from the domain $\mathcal{H}_{\psi,\omega}^{\Phi}$ onto $L^{2}_{\psi,\omega}(\mathbb{R})$. The unique solution is given by the weighted convolution:
$u(t)=\frac{1}{\omega(t)}\int_{-\infty}^{t}\mathcal{G}_{\lambda}(\psi(t)-\psi(\tau))\,f(\tau)\,\omega(\tau)\,\psi^{\prime}(\tau)\,d\tau,$ (34)
where the Green’s function $\mathcal{G}_{\lambda}$ is determined by the inverse Fourier transform of the resonant symbol:
$\widehat{\mathcal{G}_{\lambda}}(\xi)=\frac{1}{i\xi\Phi(i\xi)+\lambda}.$ (35)
Proof.
Applying the Weighted Fourier Transform $\mathcal{F}_{\psi,\omega}$ to Eq. (32) and utilizing Theorem 3.14 (with $n=1$), the differential equation converts into an algebraic one in the frequency domain:
$(i\xi\Phi(i\xi)+\lambda)\mathcal{F}_{\psi,\omega}\{u\}(\xi)=\mathcal{F}_{\psi,\omega}\{f\}(\xi).$
Under condition (33), the multiplier $m(\xi)=(i\xi\Phi(i\xi)+\lambda)^{-1}$ belongs to $L^{\infty}(\mathbb{R})$. Since the Weighted Fourier Transform is a unitary isomorphism between $L^{2}_{\psi,\omega}$ and $L^{2}$, the solution is uniquely determined in the spectral domain. The explicit time-domain representation (34) follows directly from the inverse structure $\mathcal{F}_{\psi,\omega}^{-1}=M_{\omega}^{-1}Q_{\psi}\mathcal{F}^{-1}Q_{\psi}^{-1}M_{\omega}$, which maps the product of symbols back to the weighted convolution integral. ∎
4.2 Application: Weighted Distributed Order and Ultra-Slow Transitions
A significant advantage of the Weyl-Sonine framework is its ability to encapsulate operators of distributed order. These operators model complex multi-scale memory effects and "ultra-slow" diffusion processes where the mean squared displacement grows logarithmically rather than as a power law.
Our framework allows for a natural generalization of distributed order operators to time-varying domains via the weights $(\psi,\omega)$.
Definition 4.3 (Weighted Distributed Order Operator).
Let $b:[0,1]\to[0,\infty)$ be a probability density function representing the distribution of differentiation orders. We define the Weighted Distributed Order Operator, $\mathbb{D}_{\psi,\omega}^{b}$, as the integral superposition of the weighted fractional derivatives:
$\mathbb{D}_{\psi,\omega}^{b}u(t):=\int_{0}^{1}b(\alpha)\,\mathfrak{D}_{\psi,\omega}^{\alpha}u(t)\,d\alpha.$ (36)
It is straightforward to show that $\mathbb{D}_{\psi,\omega}^{b}$ belongs to the Sonine class defined in Section 3. Specifically, it corresponds to the effective kernel $\kappa_{\text{eff}}(t)=\int_{0}^{1}\frac{b(\alpha)}{\Gamma(1-\alpha)}t^{-\alpha}\,d\alpha$.
Crucially, since $s^{\alpha}$ is a Bernstein function for $\alpha\in(0,1)$ and $b(\alpha)\geq 0$, the effective spectral symbol
$\Phi_{b}(s)=\int_{0}^{1}b(\alpha)s^{\alpha}\,d\alpha$
is itself a Complete Bernstein Function. This guarantees that the distributed order operator preserves the positivity and monotonicity properties established in our main theorems.
4.2.1 Example: Retarded Aging via Logarithmic Diffusion
Consider the canonical case of a uniform distribution of orders, $b(\alpha)=1$. The spectral symbol is computed directly:
$\Phi_{\text{uni}}(i\xi)=\int_{0}^{1}(i\xi)^{\alpha}\,d\alpha=\frac{i\xi-1}{\ln(i\xi)}.$ (37)
Applying Theorem 3.14, the fundamental solution (Green’s function) for the associated relaxation problem is governed by the multiplier:
$\widehat{\mathcal{G}}(\xi)=\frac{1}{\Phi_{\text{uni}}(i\xi)+\lambda}=\frac{\ln(i\xi)}{i\xi-1+\lambda\ln(i\xi)}.$ (38)
In classical invariant media, this symbol corresponds to ultra-slow logarithmic diffusion (see Chechkin et al. [12]). In our weighted framework, this describes a process of retarded aging: the system relaxes logarithmically with respect to the subjective internal time $\psi(t)$, modulated by the changing density $\omega(t)$. This model is particularly relevant for biological growth or materials with evolving microstructure where the "clock" slows down over time.
5 Conclusions and Perspectives
In this work, we have successfully constructed a unified mathematical framework that bridges the gap between General Fractional Calculus and the analysis of heterogeneous, aging media. By defining the Weighted Weyl-Sonine operators, we have extended the scope of non-local theory to the entire real line $\mathbb{R}$, allowing for memory kernels of arbitrary nature—governed by Complete Bernstein Functions—while simultaneously incorporating time-dependent material properties via the admissible structure pair $(\psi,\omega)$.
The core contribution of this paper is the establishment of the Generalized Spectral Mapping Theorem. We have demonstrated that, under the appropriate conjugation, the Weighted Fourier Transform acts as a unitary diagonalization map for these complex operators. This result is of significant practical value: it reduces the solution of variable-coefficient integro-differential equations to algebraic manipulations involving the spectral symbol $\Phi(i\xi)$.
The versatility of this approach was illustrated through the solvability of generalized relaxation models. We derived explicit Green’s functions for arbitrary admissible symbols, ensuring the preservation of positivity in the resulting dynamics. In particular, we demonstrated that our framework naturally encompasses Weighted Distributed Order Operators, proving that phenomena of ultra-slow diffusion and retarded aging can be rigorously analyzed and solved using the distributed spectral symbol $\Phi_{b}(\xi)$.
Furthermore, the derivation of the Marchaud-type representation (Theorem 3.17) provides a crucial link between the algebraic operator definition and the analytical perspective of finite differences and Lévy measures. This representation is theoretically significant as it relaxes the regularity requirements on the function space, and practically important as it opens the door to robust numerical implementations based on hypersingular quadratures.
Looking forward, several promising research avenues emerge from this study:
  • Weighted Sonine-Sobolev Spaces: As hinted in our spectral analysis, a rigorous treatment of the functional spaces $\mathcal{H}_{\psi,\omega}^{\Phi}$ is necessary. Future work will address embedding theorems and trace properties for these domains, providing the functional analytic bedrock for non-linear problems.
  • Inverse Problems: The explicit relation between the spectral symbol $\Phi$ and the Green’s function suggests the possibility of reconstructing the memory kernel or the aging weight $\omega(t)$ from spectral data. This is particularly relevant for identifying material properties in experimental rheology.
  • Numerical Schemes: The Marchaud representation derived here lends itself naturally to the development of fast numerical schemes, avoiding the singularity issues often found in direct Riemann-Liouville discretizations of aging problems.
We believe this framework provides the necessary mathematical tools to model the next generation of complex systems where memory, heterogeneity, and aging coexist.
Appendix A Well-posedness of Convolution Integrals
In this appendix, we provide a rigorous justification for the convergence of the Weighted Weyl-Sonine integrals defined in Section 3, ensuring they are well-defined for kernels with both local singularities and tempered growth at infinity.
Proposition A.1.
Let $k\in L_{loc}^{1}(\mathbb{R})$ be a kernel satisfying the polynomial growth condition $|k(t)|\leq C(1+|t|)^{M}$ for large $|t|$. Let $u\in\mathcal{S}_{\psi,\omega}$ be a test function in the Weighted Schwartz Space. Then, the convolution integral defining the operator converges absolutely for every $t\in\mathbb{R}$.
Proof.
By the structural factorization theorem, the operator is unitarily equivalent to the convolution (kv)(y)(k*v)(y) on the rectified line, where $v\in\mathcal{S}(\mathbb{R})$ is the conjugated test function. To prove convergence, we split the integral into a local singularity region and a distal tail region:
$|(k*v)(y)|\leq\int_{|z|\leq 1}|k(z)||v(y-z)|\,dz+\int_{|z|>1}|k(z)||v(y-z)|\,dz.$
1. Local Singularities: Since $v\in\mathcal{S}(\mathbb{R})$, it is bounded ($v\in L^{\infty}$). For the local part, we rely on the local integrability of $k$:
$\int_{|z|\leq 1}|k(z)||v(y-z)|\,dz\leq\|v\|_{\infty}\int_{-1}^{1}|k(z)|\,dz<\infty.$
This ensures that weak singularities (e.g., $z^{-\alpha}$ with $0<\alpha<1$) do not prevent convergence.
2. Growth at Infinity: For the tail $|z|>1$, we exploit the rapid decay of the Schwartz function $v$. For any $N>0$, there exists a constant $C_{N}$ such that $|v(y-z)|\leq C_{N}(1+|y-z|)^{-N}$. We invoke the classic Peetre’s inequality, which states that for any $s\in\mathbb{R}$, $(1+|a+b|)^{s}\leq(1+|a|)^{s}(1+|b|)^{|s|}$. Applying this with exponent N-N, we have:
$(1+|y-z|)^{-N}\leq(1+|y|)^{N}(1+|z|)^{-N}.$
Substituting this bound into the integral and using the polynomial growth of $k$:
$\int_{|z|>1}|k(z)||v(y-z)|\,dz\leq C\cdot C_{N}(1+|y|)^{N}\int_{|z|>1}|z|^{M}(1+|z|)^{-N}\,dz.$
Choosing $N>M+1$ ensures that the integral on the right-hand side converges. Thus, the convolution is well-defined pointwise and inherits the polynomial growth of the kernel. ∎
Appendix B Extension to Weighted Distributions
In this section, we justify the application of the Weighted Fourier Transform $\mathcal{F}_{\psi,\omega}$ to the singular kernels involved in our operator definitions. This extension relies on the topological duality established in our previous work regarding the space $\mathcal{S}_{\psi,\omega}$.
B.1 The Weighted Dual Space
We define the space of Weighted Tempered Distributions, denoted by $\mathcal{S}^{\prime}_{\psi,\omega}$, as the topological dual of the Weighted Schwartz Space $\mathcal{S}_{\psi,\omega}$. A linear functional $T$ belongs to $\mathcal{S}^{\prime}_{\psi,\omega}$ if and only if its conjugated version $\widetilde{T}=\mathcal{C}T$ defines a classical tempered distribution in $\mathcal{S}^{\prime}(\mathbb{R})$, where $\mathcal{C}=Q_{\psi}^{-1}M_{\omega}$ is the canonical isomorphism.
B.2 Definition via Parseval’s Identity
Since $\mathcal{F}_{\psi,\omega}$ is a unitary isomorphism on the test space $\mathcal{S}_{\psi,\omega}$, we can define its action on distributions via the standard duality method (Parseval’s identity).
Definition B.1 (Distributional Weighted Fourier Transform).
Let $T\in\mathcal{S}^{\prime}_{\psi,\omega}$ be a weighted tempered distribution. We define its Weighted Fourier Transform, denoted by $\widehat{T}=\mathcal{F}_{\psi,\omega}[T]$, as the unique distribution satisfying:
$\langle\widehat{T},\varphi\rangle:=\langle T,\mathcal{F}_{\psi,\omega}[\varphi]\rangle,\quad\forall\varphi\in\mathcal{S}_{\psi,\omega}.$ (39)
Remark B.2.
This definition is consistent with the Plancherel isometry on $L^{2}_{\psi,\omega}$. It ensures that for any convolution kernel $k$ (viewed as a distribution), the spectral mapping property holds rigorously:
$\mathcal{F}_{\psi,\omega}\left[\mathfrak{D}_{\psi,\omega}^{(k)}u\right]=\mathcal{F}_{\psi,\omega}[k]\cdot\mathcal{F}_{\psi,\omega}[u]$
in the sense of distributions, where the product is understood as the multiplication of a distribution by a smooth function (the spectral symbol) or via the exchange formula.